# Inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities

### v1

@@ 4.2 Existence of solutions for subcritical superscaled problems
Proofof Theorem 1.1. Lemma 4.1 ensures that $\Phi$ satisfies the (PS) condition. For $G(x,t)=|x|^{-\eta}H(t)$, let us denote $\tilde{F},\tilde{G}:E^{q}_{b}\to\mathbb{R}$ by
$\tilde{F}(u)=\int_{\mathbb{R}^{N}}F(x,u)dx=\frac{\lambda}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\int_{\mathbb{R}^{N}}\frac{H(u)}{|x|^{\eta}}dx\quad\text{and}\quad\tilde{G}(u)=\int_{\mathbb{R}^{N}}\frac{H(u)}{|x|^{\eta}}dx,$
and $\tilde{g}=\tilde{G}^{\prime}$. For $\ell_{i}=\ell(\eta_{i},r_{i})=r_{i}\delta+\eta-N$, $i=1,2$, due to (1.8), we verify that
$|\tilde{G}(u_{t})|$ $\leq\frac{c_{1}t^{\,\ell_{1}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}}}{|x|^{\eta}}dx+\frac{c_{2}t^{\,\ell_{2}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta}}dx,$
$|\tilde{g}(u_{t})v_{t}|$ $\leq c_{1}t^{\,\ell_{1}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}-1}|v|}{|x|^{\eta}}dx+{c_{2}t^{\,\ell_{2}}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}-1}|v|}{|x|^{\eta}}dx\leq c_{1}t^{\,\ell_{1}}\|u\|_{L^{r_{1}}_{\eta}}^{r_{1}-1}\|v\|_{L^{r_{1}}_{\eta}}+c_{2}t^{\,\ell_{2}}\|u\|_{L^{r_{2}}_{\eta}}^{r_{2}-1}\|v\|_{L^{r_{2}}_{\eta}}$
@@
$\tilde{F}(u_{t})\geq\tilde{G}(u_{t})\geq\frac{ct^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$
Using the embeddings of $E^{q}_{b}$ in $L^{r_{i}}_{\eta}(\mathbb{R}^{N})$ and that $\ell<{\ell(\eta,r)},\ell_{i}$ (since we considered superscaled pairs), $i=1,2$, we obtain $\tilde{g}(u_{t})v_{t}=o(t^{\,\ell})\|v\|$ and $\tilde{G}(u_{t})=o(t^{\,\ell})$ as $t\to 0^{+}$}, uniformly in $u$ on bounded subsets, and
@@
$\Phi(u_{t})=I(u_{t})-\lambda J(u_{t})-\tilde{G}(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k+1}}+o(1)\right)\quad\text{as}\,\,t\to 0^{+},$
uniformly on $u\in\mathcal{M}$. Thus, since $\lambda<\lambda_{k+1}$, for small $\rho>0$ we see that $\Phi$ satisfies (4.18) on $B_{0}$. On the other hand, for $u\in A_{0}$ we have $\tilde{\Psi}(u)\leq\lambda_{k}\leq\lambda$. As $\tilde{G}\geq 0$, for all $u\in A_{0}$ and $t\in[0,\rho]$ we get
$\Phi(u_{t})=I(u_{t})-\lambda J(u_{t})-\tilde{G}(u_{t})\leq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k}}\right)\leq 0.$
@@
$\varphi_{u}(t)$ $=$ $I(u_{t})-\lambda J(u_{t})-\tilde{G}(u_{t})=t^{\,\ell}\left(1-{\lambda}J(u)\right)-\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx$
@@
$\tilde{G}(u_{t})=\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx=o(t^{\,\ell}),\quad\text{as}\,\,\,t\to 0^{+},$
@@
$\Phi(u_{t})$ $={t^{\,\ell}}I(u)-\lambda t^{\,\ell}J(u)-\tilde{G}(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda^{+}}{\lambda_{k_{0}}}+o(1)\right),\quad\text{as}\,\,\,t\to 0^{+}.$
@@ 4.3 Existence and multiplicity for critical problems
In the proof of our results, we need some estimates on the norm of $u$ in the weighted Lebesgue spaces $L^{r}_{\eta}$, and we present it in the next lemmas.
@@
$\frac{1}{p}=\frac{\theta}{r}+\frac{1-\theta}{\tilde{r}},\qquad\frac{a}{p}=\frac{\theta\eta}{r}+\frac{(1-\theta)\tilde{\eta}}{\tilde{r}}$. (4.24)
These identities give
@@
Case (ii): $\eta<a$. The argument is completely analogous: now select $\tilde{\eta}>a$, sufficiently close to $a$ so $b<\tilde{\eta}<2$, $0<\theta(\tilde{\eta})<1$ and $2^{*}_{b,q,\tilde{\eta}}<\tilde{r}<2^{*}_{\tilde{\eta}}$.
@@
$\delta\left(\frac{p+(a-\tilde{\eta})/\delta}{\tilde{r}}-1\right)=\frac{(a-\tilde{\eta})}{p(\eta-a)+(a-\tilde{\eta})(p-r)}\,(\ell(\eta,r)-\ell\,)=:K(\tilde{\eta})\big(\ell(\eta,r)-\ell\,\big).$
@@
Finally, using the interpolation inequality with distinct weights (see e.g. [7, Inequality (2.15)]), the relation (4.24) yields (4.23). ∎
@@
The proof of Theorem 1.3 and Theorem 1.5 will be based on the following
@@
The next abstract result is a direct corollary of the preceding one and will be used in the proofs of Theorem 1.10.
@@ 4.4 Problems with subscaled nonlinearities
We now focus on nonlinearities $f$ that display at least a subscaled pair $(r,\eta)$. For the proofs of Theorems 1.6-1.9, we make use of an abstract result in critical point theory (see [25, Proposition 3.36]), which we state here for reader’s convenience.
@@
Proofof Theorem 1.6. To proceed, let us recall the notations given in Definition 1 and (1.7). Since $(\eta,r)$ is a subscaled pair, we see that the potential operator $\tilde{f}$ induced by the nonlinearity satisfy
@@
is coercive (see [23, Lemma 2.15] for a proof in a more general setting). Therefore, for any $c\in\mathbb{R}$, $(PS)_{c}$ sequences are always bounded. The compactness of the embedding $E^{q}_{b}\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ implies that $\Phi$ satisfies the $(PS)_{c}$ conditions for all levels $c\in\mathbb{R}$, as a consequence of [23, Proposition 2.3].
@@
So, let us see the case $r_{1}=2^{*}_{\eta_{1}}$. Let $(u_{n})$ be a $(PS)_{c}$ sequence, for some $c\in\mathbb{R}$. Since $(u_{n})$ is bounded, let $u\in E^{q}_{b}$ be its weak limit, up to a subsequence. Then, we follow the same steps of the proof of Lemma 4.3 until (4.13). In this case here, we get
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx=-\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+o(1).$
This gives that $(u_{n})$ converges strongly to $u$ both in $E^{q}_{b}$ and $L^{2^{*}_{\eta_{1}}}_{\eta_{1}}(\mathbb{R}^{N})$. Since $\Phi$} is bounded from below in $E^{q}_{b}$, as in the proof of the previous theorem, the levels $c_{k}$ in (4.33) are bounded from below. We just need to check that $c_{k}<0$ for all $k\geq 1$. Notice that now we have, for any $u\in\mathcal{M}$,
@@
Proofof Theorem 1.8. The associated functional, given in (1.6), is in this setting given by
@@
As in the proof of Theorem 1.6, the potential operator $\tilde{f}$ satisfy
uniformly in $u$ on bounded sets of $E^{q}_{b}$, for all $v\in E^{q}_{b}$. So, the $(PS)_{c}$ condition holds for all levels $c\in\mathbb{R}$ provided $\lambda$ is not an eigenvalue. This is proved in a more general framework in [23, Lemma 2.26]. We also note that it suffices to consider the case $\lambda>0$, since when $\lambda\leq 0$ the functional is coercive and the argument follows exactly as in the proof of Theorem 1.6. Moreover, if we consider $c_{k}$ as in (4.33), the estimate $c_{k}<0$ is also exactly the same, since here we have, for all $u\in\mathcal{M}$,
@@ 5 Radial improvements
This section is devoted to exploring the improvements that arise when working in the radial subspace $E^{q}_{b,\mathrm{rad}}\subset E^{q}_{b}$. We begin by showing that the embeddings in Lemma 2.2 admit a strictly larger range of exponents in the radial setting when $0\leq\eta<b$. In contrast, when $\eta\geq b$, no improvement is obtained, and the embedding ranges coincide with those of the nonradial case. In particular, for $N=2$ we prove that one can work with singular weights $|x|^{-\eta}$ even when $\eta<b$, a feature that is not available in the nonradial setting because of the restrictions in Lemma 2.2. The main tool leading to these improved ranges is provided by [22, Theorem 18 and Proposition 26], where refined versions of the Caffarelli–Kohn–Nirenberg inequalities are established for radial smooth functions. After proving the radial embeddings, we list some existence results for radial solutions of (1.4) and compare them with those obtained in the nonradial case.
For $N\geq 2$ and $0\leq\eta<b<2<q$, we denote
$2^{\text{rad}}_{b,q,\eta}:=\frac{q(2N-2-\eta)+2(b-\eta)}{2N-2-b}.$
@@
Suppose $N\geq 2$, let $0<b<2<q<2^{*}_{b}$ and $0\leq\eta<b$. Then we have $E^{q}_{b,rad}\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ for
@@
and for $\theta\in\left(0,1\right]$ such that $\frac{1}{r}=\theta\frac{N-2}{2(N-\eta)}+(1-\theta)\frac{N-b}{q(N-\eta)}$ it holds
@@
We are going to apply CKN inequality again, for $\gamma=-\eta/r$. From the proof of Lemma 2.2, we know that when $N\geq 3$, $\sigma=-\eta/(r\theta)+(1-\theta)b/(q\theta)\leq 0$ for all $\theta\in[\theta_{1},1]$ and $\sigma\geq-1$ if $\theta=1$, since $\eta\leq 2$. Then (2.2) holds for any $u\in C_{c}^{\infty}(\mathbb{R}^{N})$, in particular for radial functions.
On the other hand, for $\theta\in(0,\theta_{1})$ if $N\geq 3$, or $\theta\in(0,1)$ if $N=2$, we have $\sigma>0$ so that (2.4) is not satisfied. Nevertheless, for radial functions we can apply [22, Theorem 18] which guarantees that (2.2) still holds if we assume (2.1) with $\theta>0$, (2.3) and consider $\sigma\in(0,N-1]$. We verify that, for $\theta\in(0,1)$, one has
@@
Remark 5.1.
Notice that if $\eta<b$ and $N\geq 3$, then from the definitions of $2^{*}_{b,q,\eta}$ and $2^{\mathrm{rad}}_{b,q,\eta}$ we obtain
$\frac{q(N-\eta)}{N-b}\;<\;2^{\mathrm{rad}}_{b,q,\eta}\;<\;2^{*}_{b,q,\eta},$
so the admissible range of exponents is strictly larger in the radial setting. Furthermore, the exponent $2^{\mathrm{rad}}_{b,q,\eta}$ is well defined when $N=2$ and $\eta<b$, whereas $2^{*}_{b,q,\eta}$ is not defined in this regime. Consequently, for $N=2$ and $\eta<b$ we obtain an entirely new interval of Sobolev exponents, allowing us to treat the problem (1.4) in this setting.
Remark 5.2.
This last lemma becomes particularly interesting when considering the case $\eta=0$, that is, embeddings of $E^{q}_{b}$ into the non-weighted Lebesgue spaces $L^{r}$. Notice that, in the nonradial case (Lemma 2.2), we cannot take $\eta=0$ when $N=2$, and if $N\geq 3$ the admissible interval reduces to the single point $r=2^{*}$. As a consequence, no compact embeddings are available. The radial case presents a better situation: in this setting, there exist nontrivial intervals for $r$ in any dimension $N\geq 2$, and the embeddings are compact except at the endpoints.
@@
All the existence and multiplicity results stated in Theorems 1.11.10, which were formulated under the admissibility condition of Definition 1 and rely only on the weighted Sobolev embeddings from Lemma 2.2, remain valid, with exactly the same proofs, when restricted to the radial subspace $E^{q}_{b,\mathrm{rad}}$, provided that the admissibility requirement on $(\eta,r)$ is replaced by the corresponding radial embedding ranges from Lemma 5.1.
For $\eta\geq b$, the radial and nonradial admissible intervals coincide, so the statements of all theorems remain unchanged. When $0\leq\eta<b$ and $N\geq 3$, however, the radial lower endpoint improves from $2^{*}_{b,q,\eta}$ to $2^{\mathrm{rad}}_{b,q,\eta}$, thereby enlarging the class of superscaled admissible pairs. Consequently, the hypotheses of Theorems 1.11.5 and 1.7 may be relaxed in the radial setting. The remaining results either involve subscaled pairs, which never occur when $\eta<b$, or require critical exponents, and therefore do not benefit from the radial improvement.
Finally, when $N=2$ and $0\leq\eta<b$, Lemma 5.1 provides a genuinely new admissible interval [ $2^{\mathrm{rad}}_{b,q,\eta},\infty)$ for $r$, provided $(\eta,r)$ is superscaled. Thus, the results in Theorems 1.1, 1.2 and 1.7 admit radial analogues even in two dimensions, where such pairs did not exist in the nonradial case.

### v2

@@ 4.2 Existence of solutions for subcritical superscaled problems
Proofof Theorem 1.1. Lemma 4.1 ensures that $\Phi$ satisfies the $(PS)$ condition. Let us denote $\tilde{F},\tilde{H}:E^{q}_{b}\to\mathbb{R}$ by
$\tilde{F}(u)=\int_{\mathbb{R}^{N}}F(x,u)dx=\frac{\lambda}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\int_{\mathbb{R}^{N}}\frac{H(u)}{|x|^{\eta}}dx\quad\text{and}\quad\tilde{H}(u)=\int_{\mathbb{R}^{N}}\frac{H(u)}{|x|^{\eta}}dx,$
and $\tilde{h}=\tilde{H}^{\prime}$. For $\ell_{i}=\ell(\eta_{i},r_{i})=r_{i}\delta+\eta-N$, $i=1,2$, due to (1.8), we verify that
$|\tilde{H}(u_{t})|$ $\leq\frac{c_{1}t^{\,\ell_{1}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}}}{|x|^{\eta}}dx+\frac{c_{2}t^{\,\ell_{2}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta}}dx,$
$|\tilde{h}(u_{t})v_{t}|$ $\leq c_{1}t^{\,\ell_{1}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}-1}|v|}{|x|^{\eta}}dx+{c_{2}t^{\,\ell_{2}}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}-1}|v|}{|x|^{\eta}}dx\leq c_{1}t^{\,\ell_{1}}\|u\|_{L^{r_{1}}_{\eta}}^{r_{1}-1}\|v\|_{L^{r_{1}}_{\eta}}+c_{2}t^{\,\ell_{2}}\|u\|_{L^{r_{2}}_{\eta}}^{r_{2}-1}\|v\|_{L^{r_{2}}_{\eta}}$
@@
$\tilde{F}(u_{t})\geq\tilde{H}(u_{t})\geq\frac{ct^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$
Using the embeddings of $E^{q}_{b}$ in $L^{r_{i}}_{\eta}(\mathbb{R}^{N})$ and that $\ell<{\ell(\eta,r)},\ell_{i}$ (since we considered superscaled pairs), $i=1,2$, we obtain $\tilde{h}(u_{t})v_{t}=o(t^{\,\ell})\|v\|$ and $\tilde{H}(u_{t})=o(t^{\,\ell})$ as $t\to 0^{+}$}, uniformly in $u$ on bounded subsets, and
@@
$\Phi(u_{t})=I(u_{t})-\lambda J(u_{t})-\tilde{H}(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k+1}}+o(1)\right)\quad\text{as}\,\,t\to 0^{+},$
uniformly on $u\in\mathcal{M}$. Thus, since $\lambda<\lambda_{k+1}$, for small $\rho>0$ we see that $\Phi$ satisfies (4.18) on $B_{0}$. On the other hand, for $u\in A_{0}$ we have $\tilde{\Psi}(u)\leq\lambda_{k}\leq\lambda$. As $\tilde{H}\geq 0$, for all $u\in A_{0}$ and $t\in[0,\rho]$ we get
$\Phi(u_{t})=I(u_{t})-\lambda J(u_{t})-\tilde{H}(u_{t})\leq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k}}\right)\leq 0.$
@@
$\varphi_{u}(t)$ $=$ $I(u_{t})-\lambda J(u_{t})-\tilde{H}(u_{t})=t^{\,\ell}\left(1-{\lambda}J(u)\right)-\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx$
@@
$\tilde{H}(u_{t})=\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx=o(t^{\,\ell}),\quad\text{as}\,\,\,t\to 0^{+},$
@@
$\Phi(u_{t})$ $={t^{\,\ell}}I(u)-\lambda t^{\,\ell}J(u)-\tilde{H}(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda^{+}}{\lambda_{k_{0}}}+o(1)\right),\quad\text{as}\,\,\,t\to 0^{+}.$
@@ 4.3 Existence and multiplicity for critical problems
In the proofs of our next results, we will need some technical estimates for the norm of $u$ in the weighted Lebesgue spaces $L^{r}_{\eta}$. These are the subject of the next lemmas.
@@
$\frac{1}{p}=\frac{\theta}{r}+\frac{1-\theta}{\tilde{r}}\qquad\text{and}\qquad\frac{a}{p}=\frac{\theta\eta}{r}+\frac{(1-\theta)\tilde{\eta}}{\tilde{r}}$, (4.24)
so that using the interpolation inequality with distinct weights (see e.g. [7, Inequality (2.15)]), we get (4.23). The identities in (4.24) are equivalent to
@@
Case (ii): $\eta<a$. The argument is completely analogous: now select $\tilde{\eta}>a$, sufficiently close to $a$ so that $b<\tilde{\eta}<2$, $0<\theta(\tilde{\eta})<1$ and $\max\{1,2^{*}_{b,q,\tilde{\eta}}\}<\tilde{r}<2^{*}_{\tilde{\eta}}$.
@@
${\ell-\ell(\tilde{\eta},\tilde{r})}=\frac{{\tilde{r}}}{r}\,\frac{\theta(\tilde{\eta})}{1-\theta(\tilde{\eta})}\,\big(\ell(\eta,r)-\ell\,\big).$
@@
The proofs of Theorem 1.3 and Theorem 1.5 will be based on the following
@@
The next abstract result is a direct corollary of the preceding one and will be used in the proof of Theorem 1.10.
@@ 4.4 Problems with subscaled nonlinearities
We now focus on nonlinearities $f$ that display at least a subscaled pair $(\eta,r)$. For the proofs of Theorems 1.6-1.9, we make use of an abstract result in critical point theory (see [25, Proposition 3.36]), which we state here for reader’s convenience.
@@
Proofof Theorem 1.6. To proceed, let us recall the notations given in Definition 1 and (1.7). Since $(\eta,r)$ is a subscaled pair, we see that the potential operator $\tilde{f}$ induced by the nonlinearity $f(x,t)=|x|^{-\eta}|t|^{r-2}t$ satisfies
@@
is coercive (see [23, Lemma 2.15] for a proof in a more general setting). Therefore, for any $c\in\mathbb{R}$, $(PS)_{c}$ sequences are always bounded. The compactness of the embedding $E^{q}_{b}\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ (see Corollary 2.4) implies that $\Phi$ satisfies the $(PS)_{c}$ conditions for all levels $c\in\mathbb{R}$, as a consequence of [23, Proposition 2.3].
@@
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx=-\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{r_{1}}}{|x|^{\eta_{1}}}dx+o(1).$
This gives that $(u_{n})$ converges strongly to $u$ both in $E^{q}_{b}$ and $L^{r_{1}}_{\eta_{1}}(\mathbb{R}^{N})$. Since $\Phi$} is bounded from below in $E^{q}_{b}$, as in the proof of the previous theorem, the levels $c_{k}$ in (4.33) are bounded from below. We just need to check that $c_{k}<0$ for all $k\geq 1$. Notice that now we have, for any $u\in\mathcal{M}$,
@@
Proofof Theorem 1.8. The associated functional, defined in (1.6), is in this setting given by
@@
Here $f(x,t)=\lambda|x|^{-a}|t|^{p-2}t+h(x,t)$, for $h(x,t)=|x|^{-\eta}|t|^{r-2}t$, and the associated potential operator $\tilde{f}=\lambda\mathcal{B}+\tilde{h}$ is compact (due to Corollary 2.3 and 2.4) with
$\tilde{h}(u_{t})v_{t}=\int_{\mathbb{R}^{N}}\frac{|u_{t}|^{r-2}u_{t}v_{t}}{{|x|^{\eta}}}dx=o(t^{\,\ell})\|v\|\quad\mbox{as}\,\,t\to\infty,$
uniformly in $u$ on bounded sets of $E^{q}_{b}$, for all $v\in E^{q}_{b}$. So, the $(PS)$ condition holds provided $\lambda\notin\sigma(\mathcal{A},\mathcal{B})$. This is proved in a more general framework in [23, Lemma 2.26]. We also note that it suffices to consider the case $\lambda>0$, since when $\lambda\leq 0$ the functional is coercive and the argument follows exactly as in the proof of Theorem 1.6. Moreover, if we consider $c_{k}$ as in (4.33), the estimate $c_{k}<0$ is also exactly the same, since here we have, for all $u\in\mathcal{M}$,
@@ 5 Radial improvements
This section is devoted to exploring the improvements that arise when working in the radial subspace $E^{q}_{b,\mathrm{rad}}\subset E^{q}_{b}$. When $\eta>b\geq 0$, no improvement is obtained, and the embedding ranges coincide with those of the nonradial case, presented in Lemma 2.2. In contrast, as we can see in Lemma 5.1, the embeddings admit a strictly larger range of exponents in the radial setting when $0\leq\eta<b$. In particular, for $N=2$ we prove that one can work with singular weights $|x|^{-\eta}$ even when $\eta<b$, a feature that is not available in the nonradial setting. Moreover, in $E^{q}_{b,\rm rad}$ we reach compactness of the embeddings in the case $\eta=b=0$ for $N\geq 3$ and in the case $\eta=b\geq 0$ for $N=2$. Notice that in these cases Lemma 2.2 provides a nontrivial interval for continuous embedding, without compactness.
The main tool leading to these improved ranges is provided by [22, Theorem 18 and Proposition 26], where refined versions of the Caffarelli–Kohn–Nirenberg inequalities are established for radial smooth functions. After proving the radial embeddings, we list some existence results for radial solutions of (1.4) and compare them with those obtained in the nonradial case.
To denote this new interval of embeddings, for $N\geq 2$, $0\leq\eta<N$ and $0\leq b<2<q$, we consider
$2^{\text{rad}}_{b,q,\eta}:=\begin{cases}\vskip 5.69046pt\frac{q(N-\eta)}{N-b}&if\quad b\leq\eta,\\ \frac{q(2N-2-\eta)+2(b-\eta)}{2N-2-b}&if\quad 0\leq\eta<b.\end{cases}$
Notice that if $\eta\geq b$ then $2^{\mathrm{rad}}_{b,q,\eta}=2^{*}_{b,q,\eta}$. In particular, $2^{\mathrm{rad}}_{b,q,b}=2^{*}_{b,q,b}=q$. On the other hand, for $0\leq\eta<b$ we obtain
$q<\frac{q(N-\eta)}{N-b}\;<\;2^{\mathrm{rad}}_{b,q,\eta}\;<\;2^{*}_{\eta}\;\;(\;2^{\mathrm{rad}}_{b,q,\eta}\;<\;2^{*}_{b,q,\eta}\quad\text{if}\quad N\geq 3).$
We are able to prove the following:
@@
Suppose $N\geq 2$, let $0\leq b<2<q<2^{*}_{b}$ and $0\leq\eta\leq b$. Then we have $E^{q}_{b,rad}\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ for
@@
and for $\theta\in\left[0,1\right]$ such that $\frac{1}{r}=\theta\frac{N-2}{2(N-\eta)}+(1-\theta)\frac{N-b}{q(N-\eta)}$ it holds
@@
We are going to apply CKN inequality again, for $\gamma=-\eta/r$. If $\theta=0$, since $\gamma$ satisfies (2.1) we get $\eta=b$ and $r=q\ (=2^{\rm{rad}}_{b,q,b})$ and so the inequality obviously holds. Let us consider $\theta>0$. From the proof of Lemma 2.2, for $N\geq 3$ we get $\sigma=-\eta/(r\theta)+(1-\theta)b/(q\theta)\leq 0$ for all $\theta\in(0,1]$ if $\eta=b$ and for all $\theta\in[\theta_{1},1]$ when $0\leq\eta<b$. Moreover, $\sigma\geq-1$ if $\theta=1$, since $\eta\leq 2$. When $N=2$ and $\eta=b$ we get $\sigma=0$ for all $\theta\in(0,1]$. Then (2.2) holds for any $u\in C_{c}^{\infty}(\mathbb{R}^{N})$, in particular for radial functions, for $r$ related to these $\theta$.
On the other hand, when $0\leq\eta<b$, for $\theta\in(0,\theta_{1})$ if $N\geq 3$, or $\theta\in(0,1)$ if $N=2$, we have $\sigma>0$ so that (2.4) is not satisfied. Nevertheless, for radial functions we can apply [22, Theorem 18] which guarantees that (2.2) still holds if we assume (2.1) with $\theta>0$, (2.3) and consider $\sigma\in(0,N-1]$. We verify that, for $\theta\in(0,1)$, one has
@@
All the existence and multiplicity results stated in Theorems 1.11.10, which were formulated under the admissibility condition of Definition 1 and rely only on the weighted Sobolev embeddings from Lemma 2.2, remain valid, with exactly the same proofs, when restricted to the radial subspace $E^{q}_{b,\mathrm{rad}}$. In case $0\leq\eta\leq b$, some of them can be improved provided that the admissibility requirement on $(\eta,r)$ is replaced by the alternate definition:
Definition 3.
Assume ($\mathcal{H}$). We say that the pair $(\eta,r)$ is admissible if
$0\leq\eta<\frac{N+2}{2}\quad\text{and}\quad r\in(\,2^{\rm{rad}}_{b,q,\eta},\;2^{*}_{\eta}\,),\quad\text{with}\quad r>1.$
In addition, when $\eta\leq b$ we allow $r=2^{\rm{rad}}_{b,q,\eta}$, and if $\eta\leq 2$ and $N\geq 3$ we include the second endpoint $r=2^{*}_{\eta}$. The classification of these admissible pairs as scaled, subscaled and superscaled follows the same one as in Definition 1.
We use this new definition on both theorems below. The first one comprises the results of Theorems 1.1 and 1.2.
Theorem 5.2.
Assume ($\mathcal{H}$). Suppose $f(x,t)=\lambda|x|^{-a}|t|^{p-2}t+|x|^{-\eta}h(t)$, with $\lambda\in\mathbb{R}$, $0\leq\eta<N-q(N-2)/2$ and $h=H^{\prime}$ for $H:\mathbb{R}\to\mathbb{R}$ a $C^{1}$ function such that
$|h(t)|\leq c_{1}|t|^{r_{1}-1}+c_{2}|t|^{r_{2}-1}\quad\text{and}\quad c|t|^{r}\leq rH(t)\leq h(t)t,\quad\forall t\in\mathbb{R},$
for some constants $c,c_{1},c_{2}>0$, and superscaled pairs $(\eta,r_{1})$, $(\eta,r_{2})$ and $(\eta,r)$, with $2^{\rm{rad}}_{b,q,\eta}<r_{1},r_{2}<2^{*}_{\eta}$ and $q<r<2^{*}_{\eta}$. Then:
i)
If $\lambda$ is not an eigenvalue of (1.1), then problem (1.4) has a solution with positive energy.
ii)
If $\lambda\in\mathbb{R}$ then problem (1.4) has a nontrivial radial solution.
iii)
If $\lambda\in\mathbb{R}$ and $h$ is an odd function, then problem (1.4) has a sequence $(u_{k})$ of nontrivial radial solutions satisfying $\Phi(u_{k})\nearrow\infty$.
The next result describes the improvements that can be obtained in Theorems 1.3 and 1.5 and Corollary 1.4 in the radial setting.
Theorem 5.3.
Assume $N\geq 3$ and ($\mathcal{H}$). Consider $f$ as in (1.9), with $\mu\geq 0$, $\eta_{1}$ satisfying (1.10) and the superscaled pair $(\eta_{2},r_{2})$ satisfying one of the following conditions:
$\eta_{2}\in[0,\eta_{1})\quad\text{and}\quad 2^{*}_{\eta_{1}}\leq r_{2}<2^{*}_{\eta_{2}}\quad\text{or}$
$\eta_{2}\in[\eta_{1},2)\,\,(\eta_{2}=\eta_{1}\,\,\text{only if}\,\,\eta_{1}>b)\quad\text{and}\quad r_{2}<2^{*}_{\eta_{2}},\quad\text{with}\,\,r_{2}>q+(2^{*}_{\eta_{1}}-q)\frac{(b-\eta_{2})}{b-\eta_{1}}\quad\text{if}\quad\eta_{2}<b.$
The following results hold:
  • i)
    If for some $k,m\in\mathbb{N}$ it holds that
    $\lambda_{k}=\lambda_{k+1}=\dots=\lambda_{k+m-1}<\lambda_{k+m},$
    then there is a number $\delta_{k}>0$ such that problem (1.4) possesses at least $m$ distinct pairs of nontrivial radial solutions with positive energy whenever $\lambda\in\bigl(\lambda_{k}-\delta_{k},\;\lambda_{k}\bigr).$
  • ii)
    For any arbitrary $\lambda\in\mathbb{R}$ and $m\in\mathbb{N}$, there exists $\mu_{m}>0$ such that, for all $\mu>\mu_{m}$, problem (1.4) admits at least $m$ distinct pairs of nontrivial solutions with positive energy.
Another result that benefits from the radial setting is Theorem 1.7. There, the admissible pair $(\eta_{1},r_{1})$ is required to satisfy $\ell(\eta_{1},r_{1})>\ell(\eta,r)$ when $\lambda\leq 0$, or $\ell(\eta_{1},r_{1})>\ell(a,p)$ when $\lambda>0$. In both cases, the admissible range for $r_{1}$ enlarges when $0\leq\eta<b$, if Definition 3 is used in place of Definition 1. The remaining results (namely, Theorems 1.6, 1.8, 1.9, and 1.10) either involve subscaled pairs, which never occur when $\eta\leq b$, or require critical exponents, and therefore do not benefit from the radial improvement.